40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9626     Entry time: Wed Feb 12 11:57:55 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Musings on RFPDs 

[Rana, Jenne, Manasa]

We looked at the I vs. Q separation in several of the Refl PDs, while driving the PRM, while the  PRMI was locked on sidebands.

For REFL 55, we adjusted the demod phase to try to minimize the peak in the Q signal, and were only able to get it to be about 1/10th the size of the I peak.  This is not good, since it should be more like 1/100, at least.

For both REFL 11 and REFL 165, we were able to get the Q peaks to less than 1/100 of the I peak height. 

We changed the REFL55 phase from 17 to 16, and the REFL165 phase from -160.5 to -162.5. 

Since we believed that we had done a good job of setting the demod phase for REFL165, we used it to also check the balance of BS/PRM for MICH locking.  I drove the BS with an arbitrary number (0.5), which creates a peak in the I phase of REFL165, and then I put in a drive on the PRM and tweaked it around until that peak was minimized.  I came up with the same ratio as Koji had last Friday:  BS=0.5, PRM=-0.2625.  (The old ratio we were using, up until ~December when we started locking MICH with the ITMs, was BS=0.5, PRM=-0.267). 

Also, while we were locked using REFL55 I&Q, we noticed that the other REFL PDs had lots of broadband noise in their I signals, as if some noise in the REFL55 diode is being injected into the PRM, that we are then seeing in the other PDs. 

Some checks that we need to do:

* Inject a calibration line, set all the peak heights equal, and look at the noise floors of each PD.

* Use the calibration line to calibrate the PDs (especially REFL165) into meters, so that we know that it's noise is low enough to hold the PRC through the CARM offset reduction.

* Check out the state of the transmission QPDs - what is their noise, and is it good enough to use for holding the arms after we transition from green beatnote locking?  Does the whitening switching do anything?  What is the state of the whitening?

ELOG V3.1.3-